OpenStudy (anonymous):

Imma be real about this, I just need the answer. 4+4

OpenStudy (solomonzelman):

If you go 4+4+4+4+........ then it diverges (any infinite arithmetic series will diverge that is why there is no such a thing)

OpenStudy (solomonzelman):

Consider your two hands! Take your fingers EXCEPT THUMBS. Count your fingers on one hand and on the other. How many fingers altogether?

OpenStudy (kosonge):

4

OpenStudy (kosonge):

no 1

Nnesha (nnesha):

 how many circles are there :@

OpenStudy (kosonge):

lol thought it was division 8

Nnesha (nnesha):

\(\huge\color{green}{\checkmark}\)

Nnesha (nnesha):

:P

OpenStudy (solomonzelman):

I can tell you ln(4+4)=ln(8)=ln(2\(^3\))=3ln(2) and ln(2) can be approximated by a power series representation of ln(x+1)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \ln(2) =\sum_{n=0}^\infty} \frac{(-1)^n}{n+1}=1-1/2+1/3-1/4+1/5-1/6+...\)